Markus Companys common stock sold for $3.50 per share at the end of this year. The company paid a common stock dividend of $0.63 per share this year. It also provided the following data excerpts from this years financial statements Ending Beginning Balance Balance $ 35,000 $ 34,000 $ 70,000 55,000 $60,700 $ 70,000 $ 165,700 159,000 557000 $586,800 $ 67,500 $ 55,500 152,000 $ 136,800 $ 133,000 $ 133,000 $405,000 $450,000 Total liabilities and stockholders equity $ 557,000 $586,800 Cash Accounts receivable Inventory Current assets lotal assets Current liabilities Total liabilities Common stock, $1 par value Total stockholders equity This Year Sales (all on account) 825,000 478,500 346,500 Net operating income $ 194,250 $ 10,500 $ 128,625 Cost of goods sold Gross margin Interest expense Net incomeRequired: 1. What is the earnings per share? (Round your answer to 2 decimal places.) Earnings per share2. What is the price-earnings ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price-earnings ratic3. What is the dividend payout ratio and the dividend yield ratio? (Do not round intermediate calculations. Round your answer to the nearest whole percentage place. i.e., 0.1234 should be considered as 12%.) Dividend payout ratio Dividend yield ratio4, what is the return on total assets (assuming a 30% tax rate)? (Round percentage answer to 1 decimal place. i.e., 0.123 should Return on total assets
Answer #1

Answers

  • Requirement 1

A

Net Income

$128,625

B

Average Common Stock shares

$133,000

C = A/B

Earnings per share

$0.97

  • Requirement 2

A

Price per share

$3.50

B

Earnings per share

$0.97

C = A/B

Price earnings ratio

3.62

  • Requirement 3

A

Dividend per share

$0.63

B

Earnings per share

$0.97

C = (A/B) x 100

Dividend payout ratio

65%

A

Dividend per share

$0.63

B

Price per share

$3.50

C = (A/B) x 100

Dividend Yield ratio

18%

  • Requirement 4

A

Net Operating Income

$194,250

B

Beginning balance - Total Assets

$586,800

C

Ending balance - Total Assets

$557,000

D = (B+C)/2

Average Total Assets

$571,900

E = A/D

Return on Total Assets

34.0%


NOTE: If Return is to be calculated on ‘total assets’ rather than ‘average total assets’, the answer would be 34.9%